# Evaluating Mathematical Reasoning Beyond Accuracy

### v1

@@ Abstract
The leaderboard of Large Language Models (LLMs) in mathematical tasks has been continuously updated. However, the majority of evaluations focus solely on the final results, neglecting the quality of the intermediate steps. This oversight can mask underlying problems, such as logical errors or unnecessary steps in the reasoning process. To measure reasoning beyond final-answer accuracy, we introduce ReasonEval, a new methodology for evaluating the quality of reasoning steps. ReasonEval employs validity and redundancy to characterize the reasoning quality, as well as accompanying LLMs to assess them automatically. Instantiated by base models that possess strong mathematical knowledge and trained with high-quality labeled data, ReasonEval achieves state-of-the-art performance on human-labeled datasets and can accurately detect different types of errors generated by perturbation. When applied to evaluate LLMs specialized in math, we find that an increase in final-answer accuracy does not necessarily guarantee an improvement in the overall quality of the reasoning steps for challenging mathematical problems. Additionally, we observe that ReasonEval can play a significant role in data selection. We release the best-performing model, meta-evaluation script, and all evaluation results at https://github.com/GAIR-NLP/ReasonEval.
@@ 1 Introduction
Mathematical reasoning, a core cognitive skill, is crucial for resolving complex problems and making informed decisions (Hendrycks et al., 2021; Lewkowycz et al., 2022), playing a significant role in large language models (LLMs) research (Azerbayev et al., 2023; Lu et al., 2023). Given its significance, reliably evaluating mathematical reasoning in LLMs becomes crucial. Current methodologies to evaluate mathematical reasoning in LLMs focus primarily on the final result (Luo et al., 2023; Chern et al., 2023; Yu et al., 2023), neglecting the intricacies of the reasoning process. For example, the OpenLLM leaderboard,111https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard a relatively well-recognized benchmark for LLMs, uses overall accuracy to assess models’ mathematical reasoning. Despite being effective to some degree, such evaluation practice could mask underlying issues such as logical errors or unnecessary steps that compromise accuracy and efficiency of reasoning steps. In this work, we argue that a desirable evaluation criterion for mathematical reasoning encompasses not only the accuracy of the final answer but also the correctness and efficiency of each step in the reasoning process. Moreover, it is imperative that the evaluation metrics be open-source and replicable to ensure transparency and reliability.
@@
Our design philosophy stems from the fact that a correct final answer does not guarantee a flawless reasoning process (Lewkowycz et al., 2022; Uesato et al., 2022), and excessive or irrelevant reasoning steps can lead to potential errors as well as increased computational costs (Zhang et al., 2023). Once these issues go unnoticed, they can cause problems in many application scenarios. For example, in K12 mathematics education, incorrect or redundant solution steps provided by LLMs could mislead students. There have been some recent works related to the above evaluation principle. Specifically, Uesato et al. (2022); Lightman et al. (2023); Wang et al. (2023) present process reward models for mathematical reasoning,222 Uesato et al. (2022) and Lightman et al. (2023) are close-source. which focus on their utility as verifiers (Cobbe et al., 2021) to boost the accuracy (i.e., by generating many candidate solutions and selecting the one ranked highest by the verifier), leaving its underlying potential to identify reasoning errors333Identifying reasoning errors and ranking reasoning steps are different tasks. As also highlighted by Tyen et al. (2023), even state-of-the-art (SOTA) models like GPT-4 (OpenAI, 2023) still struggle to detect errors in reasoning tasks. and redundancy. Clinciu et al. (2021); Golovneva et al. (2022) rely on embedding-based methods to measure the quality of general reasoning explanation, which has limitations in handling diverse reasoning steps from various models.
@@
In response to these challenges, we propose ReasonEval, a suite comprising a new evaluation methodology with defined metrics for assessing mathematical reasoning quality and corresponding LLM-based evaluators for automated calculation. As illustrated in Figure 1, ReasonEval emphasizes the validity (i.e., the step contains no mistakes in calculation and logic) of each reasoning step, and evaluates the redundancy (i.e., the step lacks utility in solving the problem but is still valid) of steps to ensure efficiency. We explore different design options for the LLM-based evaluators and show training such a high-quality evaluator for mathematical reasoning is not a straightforward task, which calls for an LLM with strong mathematical knowledge as well as specialized datasets annotated in accordance with our evaluation criteria. With these prepared, ReasonEval achieves SOTA performance on human-labeled datasets (§4.1) and can accurately detect different types of errors generated by perturbation (§4.2).
@@ 2.2 Existing Evaluation Methodology
Instead of only using the final result as a scoring criterion, some works (Sawada et al., 2023) try to measure the reasoning quality by comparing the similarity between generated and reference solution steps (i.e., and ). Although datasets like MATH (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021) provide the ground truth solutions, the existence of diverse reasoning paths leading to the same answer renders reliance on any single one of them as a reference unreliable.
@@
This method directly asks LLMs with a well-defined prompt to generate a judgment for a given generated solution and answer (Dubois et al., 2023). This approach usually requires a very powerful LLM, often GPT-4, which may not be practical for iterative model development due to cost and transparency concerns.
@@ 3.1 Design Principle
3.1.1 Task Formulation
Following this principle, we formulate such a metric design process as a three-way classification task. Given a question and the solution steps , a label { positive, neutral, negative } will be assigned with the following probability:
where the positive label indicates the step is correct and contributes to solving the question, while the neutral label represents that the step is also correct but does not make any progress. The negative label signifies an incorrect step.
3.1.2 Scoring scheme
We can define step-level scores for each step:
Assuming we have a set of scores for a solution containing N steps, We can define solution-level scores:
3.2 Evaluator Design
3.2.1 Model Architecture
@@
Our defined evaluation method (Eq.1) can be implemented by directly prompting existing LLMs or fine-tuning LLMs using supervised dataset. To make a comprehensive study, in this work we instantiate ReasonEval by different types of LLMs, which vary in base model types (e.g., Llama2 (Touvron et al., 2023) and Mistral (Jiang et al., 2023)), model sizes (e.g., 7B, 34B), and post-training strategies (e.g., continued pretraining (Azerbayev et al., 2023) and fine-tuning).
@@
Our evaluator’s architecture is identical to the base model, except that the classification head for next-token prediction is replaced with a classification head for outputting the possibilities of each class. After normalization with the softmax layer, we get the possibility for each class.
3.2.2 Fine-tuning Data
The above task formulation allows us to utilize the existing dataset, PRM800K (Lightman et al., 2023), as training data in ReasonEval. PRM800K contains about 800,000 step-level labels over 75,000 solutions. It is collected by employing humans to label the step-by-step solutions generated by GPT-4 to MATH problems. The labels for each step are identical to those mentioned in §3.1.1.
4 Experiment
4.1 Meta-Evaluations on Human Judgement Datasets
| Invalid | Redundant | |||||||
@@ 4.1 Meta-Evaluations on Human Judgement Datasets
4.1.1 Meta-Evaluation Setup
Meta-evaluation Datasets
Inspired by Zeng et al. (2024), which propose a benchmark named Meta-Reasoning-GSM8K (MR-GSM8K), we construct a meta-evaluation dataset MR-MATH to better assess how well different evaluators can detect errors in reasoning steps. It is constructed as follows: (1) To collect the first type of errors affecting the correctness of steps, we recruit undergraduates who have a solid mathematical background to label solutions generated by Abel (Chern et al., 2023) and WizardMath (Luo et al., 2023). We collect 83 samples with incorrect steps and 76 without, pinpointing the first error location in the former. All the solutions reach the correct final answers, meaning the evaluators need to judge correctness based on the process rather than the outcome. Additionally, since the style of solutions differs from the training set of PRM800K (See examples of solutions in Appendix A.1), it tests the generalization of ReasonEval to some degree. (2) For the second type of errors affecting the efficiency of problem solving process, as they are more rarer than the first one, we sample solutions from the test set of PRM800K directly, containing 150 samples with redundant steps and 150 samples without.
@@ 4.1.1 Meta-Evaluation Setup
We compare three methods to evaluate reasoning steps automatically: embedding-based methods, prompting-based methods and step-level evaluators. For the embedding-based methods, we choose ROSCOE (Golovneva et al., 2022), a SOTA method among them. Specifically, we select the semantic alignment metrics (ROSCOE-SA) and the semantic similarity metrics (ROSCOE-SS), which do not require references solution steps. For the prompting-based methods, we select GPT-3.5-turbo and GPT-4, two mainstream generation models. We use the prompt suggested by Zeng et al. (2024) for the invalid errors and the modified version for the redundant errors. For the step-level evaluators, we select 7 base models for ReasonEval: Llama-2-7B, Mistral-7B, Llemma-7B, Llemma-34B, WizardMath-7B-V1.0 (Luo et al., 2023), WizardMath-7B-V1.1, Abel-7B-002 (Chern et al., 2023). The llemma series is continuing pertaining on math-based corpus. The Abel and WizardMath series are fine-tuning for solving mathematical problems. We also select the open-source SOTA process reward model Math-shepherd-Mistral-7B (Wang et al., 2023) to compare. The settings for these methods are in Appendix A.3 and the training details for ReasonEval are in Appendix A.2.
@@ 4.1.2 Results and Analysis
We present our main results in Table 1. ReasonEval can outperform all other methods across all error types and label levels. It is noteworthy that the identification of errors at the step level is generally more challenging than at the solution level for all methods. This suggests a higher complexity in pinpointing errors at the individual step level. Furthermore, identifying redundant errors is more intricate than invalid errors due to the inherent ambiguity involved in the former.
Figure 2: Correlation between Pass@1 for the base model on MATH and AUC for the solution-level labels.
@@
| Dataset | # Solutions | Labels | Problem Source | Labeled Method |
| Math-shepherd | correct, incorrect | MATH, GSM8K | unsupervised | |
| PRM800K | positive, neutral, negative | MATH | human |
@@
In Table 1, the Mistral-7B trained on Math-Shepherd falls behind that trained on PRM800K. We summarize the information about Math-shepherd and PRM800K in Table 3. For Math-shepherd, although having 6x training data, there is more noise in the step-level labels generated with the unsupervised method, thus harming its precision in identifying errors. And the only two classes of labels also make it limited in evaluating redundancy.
Generalization Analysis
To assess the out-of-distribution generalization of ReasonEval, we evaluate its performance on MR-GSM8K, which consists of problems from GSM8K. We present the results in Table 2. We observe that both and achieve superior performance at the step level and approach the performance of GPT-4 at the solution level, demonstrating strong generalization capabilities. For Math-shepherd-Mistral-7B, it performs best in solution-level labels but lags behind in step-level labels, suggesting that the noise in step labels from Math-Shepherd negatively impacts its ability to accurately locate error steps.
4.2 Meta-Evaluations by Perturbation
To further analyze the effect of different types of errors on ReasonEval, we follow Golovneva et al. (2022) and introduce errors in the correct reasoning steps.
@@ 4.2 Meta-Evaluations by Perturbation
We sample 184 solutions without any invalid or redundant errors from the test set of PRM800K. These solutions cover various difficulty levels and subjects. We introduce 6 types of errors to one of the intermediate steps. In Table 4 we summarize the formats of the errors. All solutions are scored by ReasonEval and Math-shepherd-Mistral-7B.
Figure 3: Box-and-whisker plots of scores when meeting different types of errors at the step level. The boundaries of the whiskers are based on the 1.5 interquartile range.
@@ Results
Figure 3 presents the results. We observe the scores for the first error step (step-level scores). We observe that: (1) The validity score exhibits a significant decrease while the redundancy score remains unchanged when meeting the logical and calculation errors (random number, random operation, swap steps, remove step). (2) For the repeated step, the validity score remains high and the redundancy score increases significantly, consistent with the fact that the repetition introduces redundancy to the overall solution without impacting its validity. (3) In the case of swapping solutions, where the solution is correct but not for the given problem, the validity score decreases and the redundancy score increases. It indicates ReasonEval treats the problem and the solution as a whole, rather than assessing the solution in isolation. (4) For Math-shepherd-Mistral-7B, it shows insensitivity to the errors causing redundancy like repetition step. In light of these observations, we conclude that ReasonEval can detect different types of errors, effectively distinguishing between those that affect the solution’s validity and those that introduce redundancy. We present other analysis and the results of solution-level scores in Appendix A.4.

### v2

@@ Abstract
The leaderboard of Large Language Models (LLMs) in mathematical tasks has been continuously updated. However, the majority of evaluations focus solely on the final results, neglecting the quality of the intermediate steps. This oversight can mask underlying problems, such as logical errors or unnecessary steps in the reasoning process. To measure reasoning beyond final-answer accuracy, we introduce ReasonEval, a new methodology for evaluating the quality of reasoning steps. ReasonEval employs validity and redundancy to characterize the reasoning quality, as well as accompanying LLMs to assess them automatically. We explore different design options for the LLM-based evaluators and empirically demonstrate that ReasonEval, when instantiated with base models possessing strong mathematical knowledge and trained with high-quality labeled data, consistently outperforms baseline methods in the meta-evaluation datasets. We also highlight the strong generalization capabilities of ReasonEval. By utilizing ReasonEval to evaluate LLMs specialized in math, we find that an increase in final-answer accuracy does not necessarily guarantee an improvement in the overall quality of the reasoning steps for challenging mathematical problems. Additionally, we observe that ReasonEval can play a significant role in data selection. We open-source the best-performing model, meta-evaluation script, and all evaluation results to facilitate future research.
Code — https://github.com/GAIR-NLP/ReasonEval
@@ 1 Introduction
Mathematical reasoning, a core cognitive skill, is crucial for resolving complex problems and making informed decisions (Hendrycks et al. 2021; Lewkowycz et al. 2022), playing a significant role in large language models (LLMs) research (Azerbayev et al. 2023; Lu et al. 2023). Given its significance, reliably evaluating mathematical reasoning in LLMs becomes crucial. Current methodologies to evaluate mathematical reasoning in LLMs focus primarily on the final result (Luo et al. 2023; Chern et al. 2023; Yu et al. 2023), neglecting the intricacies of the reasoning process. For example, the OpenLLM leaderboard,111https://huggingface.co/spaces/HuggingFaceH4/open˙llm˙leaderboard a relatively well-recognized benchmark for LLMs, uses overall accuracy to assess models’ mathematical reasoning. Despite being effective to some degree, such evaluation practice could mask underlying issues such as logical errors or unnecessary steps that compromise accuracy and efficiency of reasoning steps. In this work, we argue that a desirable evaluation criterion for mathematical reasoning encompasses not only the accuracy of the final answer but also the correctness and efficiency of each step in the reasoning process. Moreover, it is imperative that the evaluation metrics be open-source and replicable to ensure transparency and reliability.
@@
Our design philosophy stems from the fact that a correct final answer does not guarantee a flawless reasoning process (Lewkowycz et al. 2022; Uesato et al. 2022), and excessive or irrelevant reasoning steps can lead to potential errors as well as increased computational costs (Zhang et al. 2023). Once these issues go unnoticed, they can cause problems in many application scenarios. For example, in K12 mathematics education, incorrect or redundant solution steps provided by LLMs could mislead students. There have been some recent works related to the above evaluation principle. Specifically, Uesato et al. (2022); Lightman et al. (2023); Wang et al. (2023) present process reward models for mathematical reasoning,222 Uesato et al. (2022) and Lightman et al. (2023) are close-source. which focus on their utility as verifiers (Cobbe et al. 2021) to boost the accuracy (i.e., by generating many candidate solutions and selecting the one ranked highest by the verifier), leaving its underlying potential to identify reasoning errors333Identifying reasoning errors and ranking reasoning steps are different tasks. As also highlighted by Tyen et al. (2023), even state-of-the-art (SOTA) models like GPT-4 (OpenAI 2023) still struggle to detect errors in reasoning tasks. and redundancy. Clinciu, Eshghi, and Hastie (2021); Golovneva et al. (2022) rely on embedding-based methods to measure the quality of general reasoning explanation, which has limitations in handling diverse reasoning steps from various models.
@@
In response to these challenges, we propose ReasonEval, a suite comprising a new evaluation methodology with defined metrics for assessing mathematical reasoning quality and corresponding LLM-based evaluators for automated calculation. As illustrated in Figure 1, ReasonEval emphasizes the validity (i.e., the step contains no mistakes in calculation and logic) of each reasoning step, and evaluates the redundancy (i.e., the step lacks utility in solving the problem but is still valid) of steps to ensure efficiency. We explore different design options for the LLM-based evaluators and show training such a high-quality evaluator for mathematical reasoning is not a straightforward task, which requires an LLM with strong mathematical knowledge as well as high-quality training data. We empirically demonstrate that ReasonEval using the optimal design can consistently outperform baseline methods by a large margin in the meta-evaluation datasets (§4.2). Moreover, we highlight the strong generalization ability of ReasonEval (§4.3).
@@ 2.2 Existing Evaluation Methodology
Instead of only using the final result as a scoring criterion, some works (Sawada et al. 2023) try to measure the reasoning quality by comparing the similarity between generated and reference solution steps (i.e., and ). Although datasets like MATH (Hendrycks et al. 2021) and GSM8K (Cobbe et al. 2021) provide the ground truth solutions, the existence of diverse reasoning paths leading to the same answer renders reliance on any single one of them as a reference unreliable.
@@
This method directly asks LLMs with a well-defined prompt to generate a judgment for a given generated solution and answer (Hao et al. 2024). This approach usually requires a very powerful LLM, often GPT-4, which may not be practical for iterative model development due to cost and transparency concerns.
@@ 3.1 Design Principle
3.2 Scoring Scheme
Following this principle, we formulate such a metric design process as a classification task. We classify each step into three classes: positive, neutral, and negative. The positive label indicates that the step is correct and contributes to solving the question, the neutral label represents that the step is correct but does not make any progress, and the negative label signifies an incorrect step. Given a question and the solution steps , each step will be assigned the probability of the three classes as follows:
The validity score, concerning only the correctness of the reasoning steps, is defined as:
The redundancy score, concerning the utility of the steps given that they are valid, is defined as:
We can aggregate the step-level scores to get the solution-level scores. We use the ‘min’ and ‘max’ operations:
We describe the detailed justification of our scoring scheme in Appendix A.
3.3 Model Architecture
@@
Our defined evaluation method (Eq.1) can be implemented by directly prompting existing LLMs or fine-tuning LLMs using supervised dataset. To make a comprehensive study, in this work we instantiate ReasonEval by different types of LLMs, which vary in base model types (e.g., Llama2 (Touvron et al. 2023) and Mistral (Jiang et al. 2023)), model sizes (e.g., 7B, 34B), and post-training strategies (e.g., continued pretraining (Azerbayev et al. 2023) and fine-tuning).
@@
Our evaluator’s architecture is identical to the base model, except that the linear layer for next-token prediction is replaced with a linear layer for outputting the possibilities of each class. After normalization with the softmax layer, we get the possibility for each class.
3.4 Fine-tuning
The above task formulation allows us to utilize the existing dataset, PRM800K (Lightman et al. 2023), as training data in ReasonEval. PRM800K contains about 800,000 step-level labels over 75,000 solutions. It is collected by employing humans to label the step-by-step solutions generated by GPT-4 to MATH problems. The label categories for the reasoning steps are identical to those mentioned in §3.2. In the training phase, the loss function only includes the last token of each step, as it contains the full information about the step. We also take the last token of each step for prediction at test time. We provide the details on training and splitting solutions in Appendix B.
4 Meta Evaluation
4.1 Meta-evaluation Setup
| MR-MATH-invalid | MR-MATH-redundant | |||||||
@@ 4.1 Meta-evaluation Setup
Meta-evaluation444Meta-evaluation refers to evaluating the performance of evaluators themselves. Datasets
Inspired by Zeng et al. (2024), which propose a benchmark named Meta-Reasoning-GSM8K (MR-GSM8K), we construct a meta-evaluation dataset MR-MATH to better assess how well different evaluators can detect errors in reasoning steps. It is constructed as follows: (1) To collect the first type of errors affecting the correctness of steps, we recruit undergraduates who have a solid mathematical background to label solutions generated by Abel (Chern et al. 2023) and WizardMath (Luo et al. 2023). We collect 83 samples with incorrect steps and 76 without, pinpointing the first error location in the former. All the solutions reach the correct final answers, meaning the evaluators need to judge correctness based on the process rather than the outcome. Additionally, since the style of solutions differs from the training set of PRM800K (See examples of solutions in Appendix C), it tests the generalization of ReasonEval to some degree. (2) For the second type of errors affecting the efficiency of problem solving process, as they are more rarer than the first one, we sample solutions from the test set of PRM800K directly, containing 150 samples with redundant steps and 150 samples without.
| MR-GSM8K-original | MR-GSM8K-reversed | ||||||||
| Solution-level | Step-level | Solution-level | Step-level | ||||||
| OOD | F1 Score | AUC | F1 Score | AUC | F1 Score | AUC | F1 Score | AUC | |
| Embedding-based Methods | |||||||||
| ROSCOE-SA | ✓ | 51.6 | 54.4 | - | - | 54.5 | 57.9 | - | - |
| ROSCOE-SS | ✓ | 49.6 | 60.1 | - | - | 49.6 | 52.1 | - | - |
| Prompting-based Methods | |||||||||
| GPT-3.5-turbo | - | 54.9 | - | 52.3 | - | 54.3 | - | 49.9 | - |
| GPT-4 | - | 81.7 | - | 69.0 | - | 72.2 | - | 52.2 | - |
| Step-level Evaluators | |||||||||
| Math-shepherd-Mistral-7b | ✗ | 86.0 | 93.9 | 73.4 | 88.5 | 77.2 | 88.0 | 59.6 | 77.9 |
[…]
| ReasonEval | ✓ | 74.1 | 90.7 | 72.8 | 91.4 | 74.4 | 86.3 | 70.5 | 90.5 |
| ReasonEval | ✓ | 81.0 | 88.1 | 73.5 | 86.8 | 76.1 | 84.1 | 69.3 | 85.0 |
@@
We compare three methods to evaluate reasoning steps automatically: embedding-based methods, prompting-based methods and step-level evaluators. For the embedding-based methods, we choose ROSCOE (Golovneva et al. 2022), a SOTA method among them. Specifically, we select the semantic alignment metrics (ROSCOE-SA) and the semantic similarity metrics (ROSCOE-SS), which do not require references solution steps. For the prompting-based methods, we select GPT-3.5-turbo and GPT-4, two mainstream generation models. We use the prompt suggested by Zeng et al. (2024) for the invalid errors and the modified version for the redundant errors. For the step-level evaluators, we select 7 base models for ReasonEval: Llama-2-7B, Mistral-7B, Llemma-7B, Llemma-34B, WizardMath-7B-V1.0 (Luo et al. 2023), WizardMath-7B-V1.1, Abel-7B-002 (Chern et al. 2023). The llemma series is continuing pertaining on math-based corpus. The Abel and WizardMath series are fine-tuning for solving mathematical problems. We also select the open-source SOTA process reward model Math-shepherd-Mistral-7B (Wang et al. 2023) to compare. The settings for these methods are in Appendix D.
@@ 4.2 Results and Analysis
We present our main results in Table 1. ReasonEval outperforms baseline methods across all error types and label levels. It is noteworthy that the identification of errors at the step level is generally more challenging than at the solution level for all methods. This suggests a higher complexity in pinpointing errors at the individual step level. Furthermore, identifying redundant errors is more intricate than invalid errors due to the inherent ambiguity involved in the former.
@@
In Table 1, the Mistral-7B trained on Math-Shepherd falls behind that trained on PRM800K. We summarize the information about Math-shepherd and PRM800K in Table 3. For Math-shepherd, although having 6x training data, there is more noise in the step-level labels generated with the unsupervised method, thus harming its precision in identifying errors. And the only two classes of labels also make it limited in evaluating redundancy. Nonetheless, the data synthesis method is promising. We leave it for future work to optimize this method to reduce noise.
| Dataset | #S | #C | Problem Source | Human Ann. |
| Math-shepherd | 440K | 2 | MATH, GSM8K | ✗ |
| PRM800K | 75K | 3 | MATH | ✓ |
4.3 Out-Of-Distribution Generalization
@@
To assess the out-of-distribution generalization of ReasonEval, we evaluate its performance on MR-GSM8K (Zeng et al. 2024). We select two distinct types of questions from the dataset. The first type of question is from the original GSM8K. The second type, termed reversed reasoning (Yu et al. 2023), involves concealing one of the inputs and requiring the computation of the missing input using the provided original answer. Both problem sets cover the full test set of GSM8K, comprising approximately 1.4K samples each. The step-by-step solutions for these questions are sampled from MetaMath-7B (Yu et al. 2023) and include human annotations labeling the correctness of each step. Since the training data of ReasonEval does not include the labeled solutions for these types of problems, this evaluation serves as a robust test of its generalization ability. All settings are identical to those described in §4.1.
@@ Results
We present the results in Table 2. We observe that both and achieve superior performance at the step level and approach the performance of GPT-4 at the solution level, demonstrating strong generalization capabilities. Additionally, ReasonEval is more robust to the question types compared to other methods. For Math-shepherd-Mistral-7B, it performs best in solution-level labels but lags behind in step-level labels, suggesting that the noise in step labels from Math-Shepherd negatively impacts its ability to accurately locate error steps.